Gerry mixes different solutions with concentrations of 25%, 40%, and 50% to get 200 liters of a 32% solution. If he uses twice as much of the 25% solution as the 40% solution, find out how may liters of each kind he uses.
Answer by ptaylor(2198)

Let x=amount of the 40% solution
Then 2x=amount of the 25% solution
And (200-3x)=amount of 50% solution
Now we know that the amount of pure stuff that exists before the mixture takes place has to equal the amount of pure stuff that exists after the mixture takes place. So our equation to solve is:
0.40x+0.25*2x+0.50(200-3x)=0.32*200 simplify
0.40x+0.50x+100-1.50x=64
-0.60x=-100+64
-0.60x=-36
x=60 liters-----amount of 40% solution
2x=120 liters--amount of 25% solution
200-3x=200-180=20 liters----amount of 50% solution
CK
60*0.40+120*0.25+0.50*20=0.32*200
24+30+10=64
64=64
Hope this helps---ptaylor